OpenStudy (anonymous):

solve x^2-4x-12=0 by completing the square.

OpenStudy (wondermath):

(x^2-4x)-12=0 (x^2-4x+16-16)-12=0 (x^2-4x+16)-28=0 (x-4)^2-28=0

OpenStudy (anonymous):

\[x^2-4x-12=0\] \[X^2-4x=12\] \[(x-2)^2=12+2^2=12+4=16\] \[(x-2)^2=16\] \[x-2=4,\text { or } x-2=-4\] \[x=4,\text { or }x=-2\]

OpenStudy (wondermath):

sorry didn't finish

OpenStudy (wondermath):

but just solve that after